OpenStudy (anonymous):

f^1(x)=3x^2-7x, f(1)=1/2 , find f(x) .

OpenStudy (anonymous):

easy peasy, whats the problem?

OpenStudy (anonymous):

do i just plug in the 1/2 for the x?

OpenStudy (anonymous):

Nope you just copy the f(x), unless you typed the question wrong.

OpenStudy (anonymous):

\[f(x)=3x^2-7x\] would be the answer

OpenStudy (anonymous):

but it gives you the f(1)= 1/2

OpenStudy (anonymous):

i just re checked the problem and the 3x^2-7x is the inverse

OpenStudy (anonymous):

It doesnt matter lol, its a trick question: You have two buckets, one holds 4 gallons, the other 1 gallon, how many buckets do you have?

OpenStudy (anonymous):

so its really f^1(x)=3x^2-7x , then it gives you f(1)=1/2, and it's telling you to find f(x)

OpenStudy (anonymous):

Unless, as I said, it wants to you to find x for f(x)=0. But you typed to find f(x).